OpenStudy (anonymous):

Estimate the circumference of the circle with the given diameter use 3.14 for π . round to the nearest unit 1. 27 in. 2. 79 in. 3. 1,809 in. 4. 151 in.

OpenStudy (anonymous):

@Mathstudent97 @mathslover

OpenStudy (anonymous):

with the given radius or diameter

OpenStudy (anonymous):

@yman @JoelSjogren @kate323 @Bribear

mathslover (mathslover):

Use this formula to calculate the circumference of the given circle with radius = 24 inches. \(\large \boxed{\textbf{Circumference} = \color{blue}{\sf 2\pi r}}\) r = radius of the circle pi = constant = 3.14

OpenStudy (anonymous):

im still abit confused here,

mathslover (mathslover):

Where are you confused Naomix? Let me know..

OpenStudy (anonymous):

i didnt understand what the next step was because i just tried to devide and tiems tables and it didnt seem to help

mathslover (mathslover):

See, you just have to put the value of r = 24 in. in the formula : Circumference = 2 * pi * r = 2 * 3.14 * 24 inches = ?